OpenStudy (anonymous):

Can anywone help me with this Algebra 2 question? Solve by completing the square : 3x^2 - 4x - 1 = 0 I could really use some help please

OpenStudy (anonymous):

Factoring?

OpenStudy (anonymous):

I think so. Im not sure if its factoring or useing the square root property

OpenStudy (anonymous):

I believe you would be factoring this equation. 3x^2 - 4x -1 = 0 (3x+1)(x-1) = 0 3x+1= 0 x-1=0 -1 -1 subtract 1 on both side +1 +1 add 1 on both side 3x = -1 then divide both side by 3 x=1 x = -1/3 Solution is -1/3, 1